# Iyama–Solberg correspondence for exact dg categories

### 04 Jan 2024 v1

@@ 1. Introduction
where $\mathrm{id}_{\Gamma}\Gamma$ denotes the injective dimension of $\Gamma\in\mathrm{mod}\Gamma$. Iyama–Solberg [15, Theorem 4.5] showed the following
@@
We refer to Definition 2.6 for the definition of the equivalence relations mentioned in Theorem 1.3.
@@
In this article, we fix a commutative ring $k$. A dg $k$-category $\mathcal{A}$ is additive if $H^{0}(\mathcal{A})$ is an additive category. Additive subcategories of additive categories are assumed to be closed under direct summands. For an additive $k$-category $\mathcal{C}$, we denote by $\mathrm{Mod}\,\mathcal{C}$ the category of $k$-linear functors $\mathcal{C}^{op}\rightarrow k\mbox{-}\mathrm{Mod}\,$. For each object $C\in\mathcal{C}$, we denote by $C^{\wedge}$ the representable right $\mathcal{C}$-module $\mathrm{Hom}_{\mathcal{C}}(?,C)\in\mathrm{Mod}\,\mathcal{C}$. We denote by $\mathrm{mod}\mathcal{C}$ the category of finitely presented right $\mathcal{C}$-modules. Dually, we denote by $\mathcal{C}\mbox{-}\mathrm{Mod}\,$ (resp. $\mathcal{C}\mbox{-}\mathrm{mod}$) the category of $k$-linear functors $\mathcal{C}\rightarrow k\mbox{-}\mathrm{Mod}\,$ (resp. the category of finitely presented left $\mathcal{C}$-modules). For a subcategory $\mathcal{D}$ of $\mathcal{C}$, a morphism $f:X\rightarrow Y$ in $\mathcal{C}$ is a left $\mathcal{D}$-approximation of $X\in\mathcal{C}$ if $Y\in\mathcal{D}$ and the induced map $\mathrm{Hom}_{\mathcal{C}}(Y,Z)\rightarrow\mathrm{Hom}_{\mathcal{C}}(X,Z)$ is surjective for each $Z\in\mathcal{D}$. The full subcategory $\mathcal{D}$ is cavariantly finite if any $X\in\mathcal{C}$ has a left $\mathcal{D}$-approximation. If $\mathcal{X}$ is a class of objects in $\mathcal{C}$, we write ${{}^{\perp}\mathcal{X}}\coloneqq\{C\in\mathcal{C}\mid\mathrm{Hom}_{\mathcal{C}}(C,X)=0\text{ for any $X\in\mathcal{X}$}\}$. For a dg $k$-category $\mathcal{P}$, we denote by $\mathrm{pretr}(\mathcal{P})$ its pretriangulated hull [3, 8, 4] and by $\mathrm{tr}(\mathcal{P})=H^{0}(\mathrm{pretr}(\mathcal{P}))$ its triangulated hull. A full additive dg subcategory $\mathcal{I}\subseteq\mathcal{P}$ is covariantly finite if $H^{0}(\mathcal{I})$ is a covariantly finite subcategory of $H^{0}(\mathcal{P})$. For an exact dg category $\mathcal{A}$, we denote by $\mathcal{D}^{b}_{dg}(\mathcal{A})$ its (bounded) dg derived category and by $\mathcal{D}^{b}(\mathcal{A})=H^{0}(\mathcal{D}^{b}_{dg}(\mathcal{A}))$ its bounded derived category [5]. We denote by $\mathrm{dgcat}$ the category of small dg $k$-categories and by $\mathrm{Hqe}$ the localization of $\mathrm{dgcat}$ with respect to the quasi-equivalences [17].
@@ 2. Iyama–Solberg correspondence
Remark 2.7.
For a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, we have that $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is stable under kernels of retractions in $\mathrm{pretr}(\mathcal{K})$. The proof is similar to that of [6, Proposition 3.10]. Via the quasi-equivalence $\mathrm{pretr}(\mathcal{K}_{\mathcal{A}})\rightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$, the canonical morphism $\mathcal{A}\rightarrow\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{A},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ identifies $\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{A},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ with $\overline{\mathcal{A}}$, cf. Lemma 2.9 (e).
@@
Let $\mathcal{A}$ be a connective $d$-minimal Auslander–Gorenstein exact dg category. Let $\mathcal{P}$ (resp. $\mathcal{I}$) be the full dg subcategory on the projectives (resp. projective-injectives) of $\mathcal{A}$. Let $\mathcal{K}$ be the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$, which are $(d+1)$-st syzygies of objects in $\mathcal{A}$. Then the following statements hold:
@@
Let $\pi^{rel}:\mathrm{per}(\Gamma_{rel})\rightarrow\mathcal{C}$ be the quotient functor. Since $\mathrm{Hom}_{\mathrm{per}\Gamma_{rel}}(\mathcal{P},\mathrm{pvd}_{e}(\Gamma_{rel}))=0$, the functor $\pi^{rel}$ restricted to the thick category generated by $\mathcal{P}$ is fully faithful. By [18, Theorem 5.42], the Higgs category $\mathcal{H}$ is equal to the intersection ${{}^{\perp_{>0}}\mathcal{P}\cap\mathcal{P}^{\perp_{>0}}}$ in $\mathcal{C}$. In particular, it is extension-closed in $\mathcal{C}$. Let $\mathcal{I}\subseteq\mathcal{P}$ be any full additive dg subcategory. We take canonical dg enhancements of the above categories, where we add the subscripts $dg$ correspondingly. Then $(\mathcal{H}_{dg},\mathcal{P}_{dg},\mathcal{I}_{dg})$ is $d$-precluster tilting for $d\geq 0$.

### 09 Feb 2024 v2

@@ 1. Introduction
where $\mathrm{id}_{\Gamma}\Gamma$ denotes the injective dimension of $\Gamma\in\mathrm{mod}\Gamma$. Iyama–Solberg [17, Theorem 4.5] showed the following
@@
We refer to Definition 2.6 for the definition of the equivalence relations mentioned in Theorem 1.3. Indeed, the framework of exact dg categories allows us to enhance the correspondence to an equivalence of $\infty$-groupoids, cf. Theorem 2.14.
@@
In this article, we fix a commutative ring $k$. A dg $k$-category $\mathcal{A}$ is additive if $H^{0}(\mathcal{A})$ is an additive category. It is connective (resp. strictly connective) if $H^{i}\mathcal{A}(A,B)=0$ (resp. $\mathcal{A}^{i}(A,B)=0$) for $i>0$ and objects $A$ and $B$ in $\mathcal{A}$. Additive subcategories of additive categories are assumed to be closed under direct summands. For an additive $k$-category $\mathcal{C}$, we denote by $\mathrm{Mod}\,\mathcal{C}$ the category of $k$-linear functors $\mathcal{C}^{op}\rightarrow k\mbox{-}\mathrm{Mod}\,$. For each object $C\in\mathcal{C}$, we denote by $C^{\wedge}$ the representable right $\mathcal{C}$-module $\mathrm{Hom}_{\mathcal{C}}(?,C)\in\mathrm{Mod}\,\mathcal{C}$. We denote by $\mathrm{mod}\mathcal{C}$ the category of finitely presented right $\mathcal{C}$-modules. Dually, we denote by $\mathcal{C}\mbox{-}\mathrm{Mod}\,$ (resp. $\mathcal{C}\mbox{-}\mathrm{mod}$) the category of $k$-linear functors $\mathcal{C}\rightarrow k\mbox{-}\mathrm{Mod}\,$ (resp. the category of finitely presented left $\mathcal{C}$-modules). For a subcategory $\mathcal{D}$ of $\mathcal{C}$, a morphism $f:X\rightarrow Y$ in $\mathcal{C}$ is a left $\mathcal{D}$-approximation of $X\in\mathcal{C}$ if $Y\in\mathcal{D}$ and the induced map $\mathrm{Hom}_{\mathcal{C}}(Y,Z)\rightarrow\mathrm{Hom}_{\mathcal{C}}(X,Z)$ is surjective for each $Z\in\mathcal{D}$. The full subcategory $\mathcal{D}$ is cavariantly finite if any $X\in\mathcal{C}$ has a left $\mathcal{D}$-approximation. If $\mathcal{X}$ is a class of objects in $\mathcal{C}$, we write ${{}^{\perp}\mathcal{X}}\coloneqq\{C\in\mathcal{C}\mid\mathrm{Hom}_{\mathcal{C}}(C,X)=0\text{ for any $X\in\mathcal{X}$}\}$. For a dg $k$-category $\mathcal{P}$, we denote by $\mathrm{pretr}(\mathcal{P})$ its pretriangulated hull [3, 9, 4] and by $\mathrm{tr}(\mathcal{P})=H^{0}(\mathrm{pretr}(\mathcal{P}))$ its triangulated hull. A full additive dg subcategory $\mathcal{I}\subseteq\mathcal{P}$ is covariantly finite if $H^{0}(\mathcal{I})$ is a covariantly finite subcategory of $H^{0}(\mathcal{P})$. For an exact dg category $\mathcal{A}$, we denote by $\mathcal{D}^{b}_{dg}(\mathcal{A})$ its (bounded) dg derived category and by $\mathcal{D}^{b}(\mathcal{A})=H^{0}(\mathcal{D}^{b}_{dg}(\mathcal{A}))$ its bounded derived category [5]. We denote by $\mathrm{dgcat}$ the category of small dg $k$-categories and by $\mathrm{Hqe}$ the localization of $\mathrm{dgcat}$ with respect to the quasi-equivalences [19].
@@ 2. Iyama–Solberg correspondence
Let $\mathcal{A}$ be a connective $d$-minimal Auslander–Gorenstein exact dg category. Let $\mathcal{P}=\mathcal{P}_{\mathcal{A}}$ (resp. $\mathcal{I}=\mathcal{I}_{\mathcal{A}}$) be the full dg subcategory on the projectives (resp. projective-injectives) of $\mathcal{A}$. Let $\mathcal{K}=\mathcal{K}_{\mathcal{A}}$ be the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$, which are $(d+1)$-st syzygies of objects in $\mathcal{A}$. Then the following statements hold:
@@
Remark 2.9.
For a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, we have that $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is stable under kernels of retractions in $\mathrm{pretr}(\mathcal{K})$. The proof is similar to that of [6, Proposition 3.10]. Via the quasi-equivalence $\mathrm{pretr}(\mathcal{K}_{\mathcal{A}})\rightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$, the canonical morphism $\mathcal{A}\rightarrow\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{\mathcal{A}},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ identifies $\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{A},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ with $\overline{\mathcal{A}}$, cf. Lemma 2.8 (e).
@@
Let $\pi^{rel}:\mathrm{per}(\Gamma_{rel})\rightarrow\mathcal{C}$ be the quotient functor. Since $\mathrm{Hom}_{\mathrm{per}\Gamma_{rel}}(\mathcal{P},\mathrm{pvd}_{e}(\Gamma_{rel}))=0$, the functor $\pi^{rel}$ restricted to the thick category generated by $\mathcal{P}$ is fully faithful. By [21, Theorem 5.42], the Higgs category $\mathcal{H}$ is equal to the intersection ${{}^{\perp_{>0}}\mathcal{P}\cap\mathcal{P}^{\perp_{>0}}}$ in $\mathcal{C}$. In particular, it is extension-closed in $\mathcal{C}$. Let $\mathcal{I}\subseteq\mathcal{P}$ be any full additive dg subcategory. We take canonical dg enhancements of the above categories, where we add the subscripts $dg$ correspondingly. Then $(\mathcal{H}_{dg},\mathcal{P}_{dg},\mathcal{I}_{dg})$ is $d$-precluster tilting for $d\geq 0$.